OpenStudy (anonymous):

a street light is at the top of a 10 foot pole. a 6 foot tall woman walks away from the pole with a speed of 5 ft/sec along a straight path. how fast is the tip of her shadow moving when she is 35 ft from the base of the pole? ***need help setting it up***

OpenStudy (anonymous):



OpenStudy (anonymous):

similar triangles should get you through this one. you know that \[\frac{y}{6}=\frac{x+y}{10}\] and so \[10y=6x+6y\] \[4y=6x\]and finally \[x=\frac{2}{3}y\]

OpenStudy (anonymous):

actually you should use \[y=\frac{3}{2}x\] since you want \[y'\] and you are told that \[x'=5\] you see that \[y'=\frac{3}{2}\times 5\]

OpenStudy (anonymous):

okay so would i need to multiply 3/2 and 5 together?

OpenStudy (anonymous):

yes that will do it.

OpenStudy (anonymous):

i did and its tellng me its wrong?